OpenStudy (marcelie):

Eliminate the parameter X=2cos t Y=1+sint

OpenStudy (cazeh):

cos(t) = x/2 cos^2(t) = x^2/4 Now do something similar with y. Then add them up. The sum is 1 Would like medal and fan. -Credits to friend in skype!

OpenStudy (marcelie):

Oops.. Suppose to be theta lol

OpenStudy (mathmale):

An alternative approach: X=2cos t Y=1+sint Solve the first equation for cos t and the second for sin t. Now square both sides of both equations. Note that \[\sin ^{2}x+\cos ^{2}x=1\]

OpenStudy (mathmale):

This fact will enable you to eliminate the trig functions and be left with a relationship in x and y only.

OpenStudy (marcelie):

(x/2)^2+ (y-1)^2=1

OpenStudy (marcelie):

This is where i get lost

OpenStudy (marcelie):

@mathmale

OpenStudy (marcelie):

@cazeh

OpenStudy (cazeh):

it is hard i cant help sry dude :d

OpenStudy (marcelie):

Lol okay

OpenStudy (dumbcow):

(x/2)^2 = x^2 / 4 \[(y-1)^2 = 1 - \frac{x^2}{4}\] convert to 1 fraction \[(y-1)^2 = \frac{4-x^2}{4}\] take sqrt \[y-1 = \frac{\pm \sqrt{4-x^2}}{2}\] \[y = 1 \pm \frac{\sqrt{4-x^2}}{2}\]

OpenStudy (marcelie):

Oh could we actually do that to eliminate the parameter?